OpenStudy (anonymous):

How many solutions does the equation −2y + 2y + 3 = 3 have? One Zero Infinitely many Three

OpenStudy (anonymous):

@callender535

OpenStudy (anonymous):

@callielovesyhuu

OpenStudy (butterflydreamer):

What do you think? First you look at your equation: -2y + 2y + 3 = 3 and you can re-write it as : 2y - 2y + 3 = 3 Simplify by collecting like terms (note 2y - 2y = 0) 0 + 3 = 3 3 = 3 Which is true... this means that infinitely, no matter what the value of y is, there will always be a solution :)

OpenStudy (butterflydreamer):

For example, if we say y = 5 and we plug it into the equation: -2(5) + 2(5) + 3 = 3 -10 + 10 + 3 = 3 3 = 3 So y = 5 is a solution Pick another random number.. say y = 100, plug it into the equation and we get: -2 (100) + 2(100) + 3 = 3 -200 + 200 + 3 = 3 3=3 Therefore y = 100 is a solution So can you see that there are infinite values of y?

OpenStudy (anonymous):

Let's solve your equation step-by-step. −2y+2y+3=3 Step 1: Simplify both sides of the equation. −2y+2y+3=3 Simplify: (Show steps) 3=3 Step 2: Subtract 3 from both sides. 3−3=3−3 0=0 Answer: All real numbers are solutions. what do you think

OpenStudy (butterflydreamer):

excellent :)